Suppose your total gross income per month is $6,000/month. Assuming that property taxes, homeowner's insurance and mortgage insurance payment total of 300/month. In addition, you have car and student loan payments that total $500 a month. If 30 year fixed rate mortgages have a current annual percentage rate of 5.5%, how much do you qualify to borrow based on the performance to income ratios given you have enough money to pay 20% down payment?
Answer #1

Debt to income ratio = 43%

Hence, permissible debt payment per month = Monthly salary x 43% = 6,000 x 43% = 2,580

Existing car & student loan payment = $ 500 / month

Hence, maximum home mortgage payment permissible per month = 2,580 - 500 = $ 2,080

----------------------

The housing expense ratio is typically required to be approximately 28% or less = 28% x 6,000 = 1,680

Hence, Monthly mortgage payment + property taxes, homeowner's insurance and mortgage insurance payment = Monthly mortgage payment + 300 = 1,680

maximum home mortgage payment permissible per month = 1,680 - 300 = 1,380

---------------

Hence, qualifying amount = min of the qualifying amounts based on the two ratios = min (2080, 1380) = $ 1,380

This is the monthly mortgage payment you are allowed. 30 year fixed rate mortgages have a current annual percentage rate of 5.5%

Hence, the amount you qualify to borrow based on the performance to income ratios given you have enough money to pay 20% down payment = - PV (Rate, Nper, PMT, FV) = - PV (5.5% / 12, 12 x 30, 1380, 0) = $  243,048